Find the selling price per pound of a coffee mixture made from 8
pounds of coffee that sells for $9.20 per pound and 12 pounds of coffee
that sells for $5.50 per pound.

Answer by Cintchr(481)

Remeber this formula .... A x N = T
$9.20 * 8 = 73.60
$5.50 *12 = 66.00
Add the two equations
Since we are finding the price, dont add the averages.
$A * 20 = 133.60
Divide by 20
$A = 6.68